If x&y are positive number such that log2x+log2y≥6 then find the least value of x+y.


 

Grade:12

3 Answers

Pratik Tibrewal
10 years ago
log2(xy) >= 6
=> xy >= 2^6

x + y >= 2*(xy)^(1/2) >= 2*(2^3) = 16

Thanks and Regards,
Pratik Tibrewal
askiitans faculty,
BTech, IITG
Krrish .
10 years ago
5
Krrish .
10 years ago
1000